OpenStudy (anonymous):

Why is abs(cos theta) never greater than 1 in the first place? (this is linear algebra)

OpenStudy (anonymous):

actually it looks like trig cosine is the first coordinate on the unit circle (at least that is one way to think of it) so it is never bigger than 1 or smaller than -1

OpenStudy (anonymous):

But I need mathematical proofs to show why it's never greater than 1.

OpenStudy (anonymous):

i guess that depends entirely on your definition of cosine

OpenStudy (anonymous):

Well, this is like the first couple chapters of linear algebra.

OpenStudy (loser66):



OpenStudy (anonymous):

I know that cos 0=1.

OpenStudy (anonymous):

So the domain is [-1, 1] for cos theta, right?

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

the "range" for sine and cosine functions is \([-1,1]\)

OpenStudy (anonymous):

you can prove from the identity \[ \begin{align*} \sin^2x+\cos^2x&=1\\ \implies\cos^2x&=1-\sin^2x \end{align*} \] since the square of a number is always positive, the right hand side is always less than "1" but must be >0 hence the square-root must also be in that range. you can prove similarly arguing using Pythagorean thm.